Find the surface integral of the vector field 𝑭(π‘₯,𝑦,𝑧)=(π‘₯,𝑦,𝑧) over the part of the paraboloid 𝑧=1βˆ’π‘₯^2βˆ’π‘¦^2 with 𝑧β‰₯0 and having normal pointing upwards.


Hint: take π‘₯ and 𝑦 as independent parameters.


1
Expert's answer
2021-11-28T18:54:51-0500

The surface S can be represented by:

"r(x, y) = x i + y j + (1 − x^ 2 − y ^2 ) k,\ −1 ≤ x ≤ 1, −1 ≤ y ≤ 1"

"r_x = i \u2212 2x k , r_y = j \u2212 2y k"

"r_x × r_y=\begin{vmatrix}
 i & j&k \\
 1 & 0&-2x\\
0 & 1&-2y
\end{vmatrix}=2xi+2yj+k"


"\iint_SF\cdot ndS=\iint F\cdot (r_x × r_y)dA=\int^1_{-1} \int^1_{-1}(2x^2+2y^2+1 − x^ 2 − y ^2)dxdy="


"=\int^1_{-1} \int^1_{-1}(x^2+y^2+1 )dxdy=\int^1_{-1} (2/3+2y^2+2)dy="


"=4\/3+4\/3+4=20\/3"